OpenStudy (anonymous):

what is the most general antiderivative of the function F(x)= ½ + ¾ x^2 -4/5 x^3

zepdrix (zepdrix):

Ok so we need to find the anti-derivative of this?\[\Large F(x) \quad=\quad \frac{1}{2}+\frac{3}{4}x^2-\frac{4}{5}x^3\]

zepdrix (zepdrix):

Are you familiar with the process for applying the `Power Rule` in reverse? Because that's what we'll be doing here.

zepdrix (zepdrix):

When we take the derivatives, we: ~Bring the exponent down as a coefficient being multiplied. ~Subtract 1 from the exponent. When we need to find an anti-derivative, we do the opposite and in the reverse order: ~Add 1 to the exponent. ~Bring the new exponent down as a coefficient being divided.

zepdrix (zepdrix):

Example: \[\Large F(x)=x^3\] f(x) represents the antiderivative of F(x):\[\Large f(x)=\frac{1}{3+1}x^{3+1}\quad=\quad \frac{1}{4}x^4\]